Question 9 : A reaction with a negative value of [Math Processing Error]
ΔG
is said to be
Reversible
Non-spontaneous
Spontaneous
At equilibrium

Question 10 : Use the equations to answer the question. [Math Processing Error]
2A+B→A2B
[Math Processing Error]
ΔH=–217.3
kJ; [Math Processing Error]
B+C→BC
[Math Processing Error]
ΔH=–867.5
kJ. What is the value of ΔH for the reaction [Math Processing Error]
2A+BC→A2B+C
?
–650.2 kJ
–1084.8 kJ
650.2 kJ
1084.8 kJ
1
Expert's answer
2017-03-13T09:25:01-0400
Solution:
Question 9: When the value ΔG is negative, the process occurs spontaneously and can proceed in the forward direction. So the correct answer: spontaneous.
Question 10: 2A+B→A2B, ΔH1 = -217.3 kJ;
B+C→BC, ΔH2 = -867.5 kJ;
2A+BC→A2B+C, ΔH = ΔH1 + (-ΔH2) = -217.3 + 867.5 = 650.2 kJ
So the correct answer: 650.2 kJ.
Answer: Question 9: spontaneous;
Question 10: 650.2 kJ.